# A Bayesian likely responder approach for the analysis of randomized controlled trials

### v1

@@ 1 Introduction
Precision medicine aims to tailor medical treatment to the individual characteristics of each patient [national2011toward]. Tracking individual differences in treatment effects has become increasingly important, as many treatments do not uniformly benefit all patients [kaiser2022heterogeneity, sundstrom2023heterogeneity]. In randomized controlled clinical trials (RCTs), there are instances where no significant differences in efficacy are observed between candidate therapies and a placebo when evaluating the group as a whole. However, this can obscure individual variations, with some participants responding well, while others derive little benefit, leading to a minimal apparent overall treatment effect [ruberg2010mean].
@@
Laska et al.[laska2022likely] introduced the Likely Responder (LR) framework to identify patients predicted, from pre-treatment characteristics, to achieve a favorable outcome under a given treatment and to conduct inference on treatment effects within this predicted LR subgroup. By first identifying these likely responders and then comparing treatment versus control within this subgroup, the LR framework seeks to provide a clinically interpretable, targeted assessment of benefit [laska2022likely, laska2020gabapentin, wastyk2023randomized, kalso2007predicting]. Our work builds on this line of research by addressing a key gap: propagating uncertainty from subgroup identification into subsequent subgroup-specific causal inference.
@@
Likely responders are identified using a prognostic score [hansen2008prognostic], which models the outcome as a function of baseline covariates within a treatment group. This score can be viewed as a prognostic balancing score (PBS), as individuals with similar PBS values are approximately “balanced” on their covariates; consequently, PBS values can be imputed counterfactually for patients in the comparator arm from their baseline characteristics. Patients are classified as likely responders when their PBS exceeds a prespecified threshold, minCond, chosen to reflect a clinically meaningful outcome level [laska2022likely, laska2020gabapentin]. Different clinical priorities may motivate different minCond values (e.g., stricter thresholds for high-risk therapies), with each choice yielding a distinct LR analysis. When the PBS model is adequate, inference proceeds by comparing outcomes between treatment and comparator within the subgroup defined by the thresholded PBS. Accordingly, treatment effect estimation under the LR framework follows a two-stage workflow: (i) construct the PBS and apply the clinical threshold to define LR membership, and (ii) estimate the subgroup-specific treatment effect within the LR set. A diagram illustrating this workflow is shown in Fig. 1.
@@
In Fig. 1, the uncertainty from both model-based prognostic score estimation in Stage 1 and treatment effect estimation in Stage 2 contribute together to the overall uncertainty of subgroup treatment effect uncertainty estimation. Both sources contribute to the sampling variability of the subgroup-specific effect estimate. However, previous applications that estimate the average treatment effect (ATE) among LRs typically ignore uncertainty from the prognostic-score estimation in subsequent downstream analyses for evaluating the subgroup-specific treatment effects [laska2020gabapentin, laska2022likely, chalkou2021two, goligher2023heterogeneous]. This oversight can lead to potentially overconfident inference on the subgroup-specific treatment effects. The need to propagate first-stage uncertainty is recognized in adjacent literatures. For example, in observational studies, investigators have used Bayesian frameworks to carry forward uncertainty from estimated propensity scores into outcome analyses [liao2020uncertainty, alvarez2021uncertain]. Our contribution adapts this uncertainty-propagation principle to a different setting: quantifying subgroup-specific treatment effects in a predictive framework, where prognostic scores are used to define clinically interpretable LR subgroups rather than to adjust for treatment-confounding.
@@
To validate our proposed methodology, we conduct simulations with nonlinear predictor-outcome relationships and both continuous and binary outcomes. We compare results between a conventional approach, which uses point estimates in the design stage followed by a generalized linear model (GLM) based on the single stratification, and our two-stage Bayesian approach, which draws posterior samples in the design stage and estimates the ATE via GLM for each posterior-induced stratification in the evaluation stage. Reproducible code for the two-stage Bayesian procedure is available at https://github.com/ad6557/Two-stage-Bayesian-ATE. We illustrate the utility of this framework through its application to the COMPILE study (Continuous Monitoring of Pooled International Trials of Convalescent Plasma for COVID-19 Hospitalized Patients) [troxel2022association]. By explicitly accounting for posterior uncertainties in data-driven subgrouping, our methodology aims to improve the assessment of subgroup-specific treatment effects and support more informed medical interventions.
@@ 2.2 Design Stage
While other Bayesian approach could be used, we employ Bayesian Additive Regression Trees (BART) to estimate $s_{i}=E\left(Y_{i}\mid T_{i}=1,X_{i}\right)$ in the “design” stage [chipman2010bart]. BART is a semi-parametric, ensemble-of-trees model that flexibly captures nonlinearity and high-order interactions via a sum of shallow regression trees with shrinkage priors. This structure accommodates many covariates and complex effect shapes while regularization mitigates overfitting. The Bayesian formulation yields posterior draws of the prognostic score, $\{s_{i}^{(k)}\}_{k=1}^{K}$, which we later use to induce posterior-driven subgroup designs. Implementation details and prior specifications are provided in Section A.2 of the Supplementary Material.
@@ 3.3 Between- and within-design uncertainty
Viewing the subgrouping $\nu$ as a missing value that can be multiply-imputed, we estimate the posterior variance of $\Delta(\nu_{0})$ using Rubin’s combining rules [rubin1987multiple].
@@ 4.2 Competing Methods
For the “design” stage that identifies LR and UR subgroups, we consider two approaches: eXtreme Gradient Boosting (XGBoost) [chen2016xgboost] and BART [chipman2010bart]. XGBoost is an optimized implementation of Gradient Boosting Machines (GBMs), a widely used ensemble learning technique [friedman2001greedy]. GBMs iteratively refine predictions by combining weak learners through gradient descent. XGBoost enhances this process with regularization, parallelized computation, efficient handling of missing values, and flexible loss functions, while further improving performance through tree depth control, subsampling, and feature selection. The details of GBM is provided in Section A1 in the Supplemental Material.
@@
The XGBoost models in the design stage were trained using the ‘xgboost’ package in R [chen2019package], and BART models were trained using the ‘BART’ package in R [sparapani2021nonparametric]. The simulation environment was coded in R v. 4.2.1. For continuous outcomes, XGBoost models were configured to minimize the squared error loss function with root mean squared error (RMSE) as the evaluation metric, while for binary outcomes, models employed logistic regression with the deviance as the evaluation metric. Hyperparameter tuning was performed via 5-fold cross-validation over a predefined grid to mitigate overfitting, adjusting learning rate (0.02, 0.05, 0.1, 0.3), maximum tree depth (3, 5, 7), and subsample ratio of features (0.8, 1.0). BART models were fitted using default hyperparameters, including a sum of 200 trees, a base prior of 0.95, and a power prior of 2, chosen to penalize excessive branch growth (see Section A2 in the Supplemental Material). Each model was trained using a single chain MCMC procedure, consisting of 500 burn-in iterations followed by 100 retained iterations for analysis.
@@ 4.3 Simulation Results
Tables 1 and S6 present the detailed results for binary and continuous outcomes. Regarding subgroup-specific treatment effect estimation, the results indicate that the “naïve” method with BART slightly outperforms the “naïve” method with XGBoost in terms of bias, variance, and MSE. Both XGBoost-trained and BART-trained models exhibit bias toward the null in subgroup-specific treatment effect estimation, though the bias in BART is slightly smaller than that in XGBoost. Additionally, the variance of the BART-trained model across 200 simulation runs is lower, indicating greater stability. As a result, the MSE is smaller for BART than for XGBoost. However, the average standard error (SE) of the ATE for both the “naïve” frequentist method and the native method with BART is smaller than that of the proposed Bayesian uncertainty-corrected method, and their 95% Confidence Interval (CI) coverages fall below 95%, indicating that the “naïve” methods underestimate the uncertainty in their estimates. In contrast, the two-stage Bayesian method maintains strong precision and accuracy while providing coverage that is close to the nominal 95%, demonstrating its superior performance in both uncertainty estimation and overall reliability.
@@ 5.1 Data Description
We apply the proposed two-stage sampling-based approach to evaluating treatment effects to data from $n=2341$ patients in the COMPILE COVID-19 clinical trial. The study evaluates the effectiveness of the COVID-19 convalescent plasma (CCP) treatment for hospitalized COVID-19 patients not on mechanical ventilation at the time of randomization [troxel2022association]. The primary outcome is a binary outcome derived from the World Health Organization (WHO) 11-point clinical scale, which assesses COVID-19 severity with values ranging from 0 (no infection) to 10 (death) [marshall2020minimal]. (see Section C in the Supplemental Material.) Specifically, we consider four binary outcomes: (1) ventilation or death, defined as an indicator for a WHO score $\geq 7$ measured at 14 ± 1 days post-randomization (the primary outcome); (2) hospitalization, ventilation, or death, defined as an indicator for a WHO score $\geq 4$ measured at 28 ± 2 days post-randomization (secondary outcome); (3) ventilation or death, defined as an indicator for a WHO score $\geq 7$ measured at 28 ± 2 days post-randomization (secondary outcome); and (4) death, defined as an indicator for a WHO score of 10 at 28 ± 2 days post-randomization (secondary outcome). Park et al.[park2022development] developed a Treatment Benefit Index to identify hospitalized COVID-19 patients who may benefit from CCP, however, their method ignores uncertainty in estimating data-driven model-based benefit groups. The treatment variable is a binary indicator (1 for CCP treatment, 0 for control). Baseline covariates used in our study are as follows:
@@ 5.2 Application Results
To estimate the estimand in each subgroup, we implemented a Monte Carlo (MC) cross-validation (CV) procedure with 100 repetitions, as in Goligher et al. [goligher2023heterogeneous]. MC–CV is employed in the application to mitigate the overfitting risk inherent to single-sample analyses, whereas the simulation study uses genuinely independent design and evaluation samples and therefore does not require cross-validation. In each MC iteration, we followed the following steps:
@@
A final predicted estimand was obtained by taking the median point estimate and confidence interval across 100 MC replications [goligher2023heterogeneous]. The reported metrics in Table S9 are as follow: 1) OR: exponentiated median of the log ORs; 2) $\mbox{SE}_{logOR}$: median of the standard errors of the log ORs; and 3) 95% CI; exponentiated median of the lower and upper bounds of the 95% confidence intervals of the log ORs.
@@ 6 Discussion
Beyond statistical considerations, the LR framework can have clinical and regulatory value. In trials with null or inconclusive average effects, LR analyses may suggest that a treatment benefits a subset of patients; in trials with positive averages, they can help characterize where responses are stronger. Within this framework, our Bayesian approach explicitly propagates uncertainty from subgroup identification to subgroup-specific inference, reducing the risk of overconfident conclusions. LR analyses also connect to enrichment in trial design. U.S. FDA guidance describes enrichment as prospectively selecting patients more likely to respond than the broader target population, thereby improving power [FDA_EnrichmentGuidance]. LR subgroups identified within an RCT offer a data-driven input to such enrichment, because potential responders are characterized using baseline features.
@@
Although the method is developed in the context of an RCT, it can be extended to more general settings. In early-phase or pilot studies where only single-arm data are available, the “design” stage can rely on treated patients alone to develop a biomarker-based score, which can then be used to define subgroups in a subsequent, independent trial. The approach is also applicable to other model-based, data-driven subgroup analyses, such as those based on stratification by heterogeneous treatment effects [thanassoulis2016individualized, lee2017coronary].
@@ A. Introduction of Tree-based Models
The XGBoost method [chen2016xgboost] is based on the concept of gradient boosting , a machine learning technique that enhances weak learners into strong ones. Friedman et al.[friedman2000additive] later adapted boosting to a forward stagewise additive model that minimizes a loss function and brought forth a highly adaptable algorithm, gradient boosting machines (GBM). Specifically, a GBM uses $m$ additive functions to predict the outcome,
@@
XGBoost further mitigates overfitting using shrinkage and column subsampling. Shrinkage scales newly added weights after each boosting step, similar to traditional gradient tree boosting. Column subsampling, inspired by the Random Forest algorithm [breiman2001random], selects a random subset of predictors for splitting at each step, enhancing model robustness.[chen2016xgboost]
@@
The BART model is a Bayesian probability model, requiring prior distribution specification. Chipman et al. [chipman2010bart] simplified this step using independence assumptions:
@@
For $T_{j}$, the prior probability that nodes at depth $d$ are non-terminal is set as $\alpha(1+d)^{-\beta}$ with base $\alpha=0.95$ and power $\beta=2$ [chipman2010bart], encouraging smaller trees. Under this distribution, tree size of $\{1,2,3,4,5\}$ has a prior probability of $\{0.05,0.55,0.28,0.09,0.03\}$, respectively.
@@
Given the prior and the observed data, posterior sampling in BART models is carried out via Markov chain Monte Carlo (MCMC) based on "Bayesian back-fitting" [hastie2000bayesian], which is an application of the "simple" Gibbs sampler. Specifically, the $j$th tree $(T_{j},M_{j})$ is fit iteratively, holding all other m1m-1 trees constant. For each iteration, given $\{T_{j}\}$ and $\{M_{j}\}$ and based on the model residual $R_{i}=Y_{i}-\sum_{j=1}^{m}g(x_{i},T_{j},M_{j})$, we draw a sample from the posterior of $\sigma^{2}$ ("Step 1"). Then, we define the $j$th partial residual response

### v2

@@ 1 Introduction
Precision medicine aims to tailor medical treatment to the individual characteristics of each patient [8]. Tracking individual differences in treatment effects has become increasingly important, as many treatments do not uniformly benefit all patients [14, 25]. In randomized controlled clinical trials (RCTs), there are instances where no significant differences in efficacy are observed between candidate therapies and a placebo when evaluating the group as a whole. However, this can obscure individual variations, with some participants responding well, while others derive little benefit, leading to a minimal apparent overall treatment effect [22].
@@
Laska et al.[17] introduced the Likely Responder (LR) framework to identify patients predicted, from pre-treatment characteristics, to achieve a favorable outcome under a given treatment and to conduct inference on treatment effects within this predicted LR subgroup. By first identifying these likely responders and then comparing treatment versus control within this subgroup, the LR framework seeks to provide a clinically interpretable, targeted assessment of benefit [17, 16, 29, 15]. Our work builds on this line of research by addressing a key gap: propagating uncertainty from subgroup identification into subsequent subgroup-specific causal inference.
@@
Likely responders are identified using a prognostic score [12], which models the outcome as a function of baseline covariates within a treatment group. This score can be viewed as a prognostic balancing score (PBS), as individuals with similar PBS values are approximately “balanced” on their covariates; consequently, PBS values can be imputed counterfactually for patients in the comparator arm from their baseline characteristics. Patients are classified as likely responders when their PBS exceeds a prespecified threshold, minCond, chosen to reflect a clinically meaningful outcome level [17, 16]. Different clinical priorities may motivate different minCond values (e.g., stricter thresholds for high-risk therapies), with each choice yielding a distinct LR analysis. When the PBS model is adequate, inference proceeds by comparing outcomes between treatment and comparator within the subgroup defined by the thresholded PBS. Accordingly, treatment effect estimation under the LR framework follows a two-stage workflow: (i) construct the PBS and apply the clinical threshold to define LR membership, and (ii) estimate the subgroup-specific treatment effect within the LR set. A diagram illustrating this workflow is shown in Fig. 1.
@@
In Fig. 1, the uncertainty from both model-based prognostic score estimation in Stage 1 and treatment effect estimation in Stage 2 contribute together to the overall uncertainty of subgroup treatment effect uncertainty estimation. Both sources contribute to the sampling variability of the subgroup-specific effect estimate. However, previous applications that estimate the average treatment effect (ATE) among LRs typically ignore uncertainty from the prognostic-score estimation in subsequent downstream analyses for evaluating the subgroup-specific treatment effects [16, 17, 4, 11]. This oversight can lead to potentially overconfident inference on the subgroup-specific treatment effects. The need to propagate first-stage uncertainty is recognized in adjacent literatures. For example, in observational studies, investigators have used Bayesian frameworks to carry forward uncertainty from estimated propensity scores into outcome analyses [19, 1]. Our contribution adapts this uncertainty-propagation principle to a different setting: quantifying subgroup-specific treatment effects in a predictive framework, where prognostic scores are used to define clinically interpretable LR subgroups rather than to adjust for treatment-confounding.
@@
To validate our proposed methodology, we conduct simulations with nonlinear predictor-outcome relationships and both continuous and binary outcomes. We compare results between a conventional approach, which uses point estimates in the design stage followed by a generalized linear model (GLM) based on the single stratification, and our two-stage Bayesian approach, which draws posterior samples in the design stage and estimates the ATE via GLM for each posterior-induced stratification in the evaluation stage. Reproducible code for the two-stage Bayesian procedure is available at https://github.com/ad6557/Two-stage-Bayesian-ATE. We illustrate the utility of this framework through its application to the COMPILE study (Continuous Monitoring of Pooled International Trials of Convalescent Plasma for COVID-19 Hospitalized Patients) [27]. By explicitly accounting for posterior uncertainties in data-driven subgrouping, our methodology aims to improve the assessment of subgroup-specific treatment effects and support more informed medical interventions.
@@ 2.2 Design Stage
While other Bayesian approach could be used, we employ Bayesian Additive Regression Trees (BART) to estimate $s_{i}=E\left(Y_{i}\mid T_{i}=1,X_{i}\right)$ in the “design” stage [7]. BART is a semi-parametric, ensemble-of-trees model that flexibly captures nonlinearity and high-order interactions via a sum of shallow regression trees with shrinkage priors. This structure accommodates many covariates and complex effect shapes while regularization mitigates overfitting. The Bayesian formulation yields posterior draws of the prognostic score, $\{s_{i}^{(k)}\}_{k=1}^{K}$, which we later use to induce posterior-driven subgroup designs. Implementation details and prior specifications are provided in Section A.2 of the Supplementary Material.
@@ 3.3 Between- and within-design uncertainty
Viewing the subgrouping $\nu$ as a missing value that can be multiply-imputed, we estimate the posterior variance of $\Delta(\nu_{0})$ using Rubin’s combining rules [23].
@@ 4.2 Competing Methods
For the “design” stage that identifies LR and UR subgroups, we consider two approaches: eXtreme Gradient Boosting (XGBoost) [5] and BART [7]. XGBoost is an optimized implementation of Gradient Boosting Machines (GBMs), a widely used ensemble learning technique [9]. GBMs iteratively refine predictions by combining weak learners through gradient descent. XGBoost enhances this process with regularization, parallelized computation, efficient handling of missing values, and flexible loss functions, while further improving performance through tree depth control, subsampling, and feature selection. The details of GBM is provided in Section A1 in the Supplemental Material.
@@
The XGBoost models in the design stage were trained using the ‘xgboost’ package in R [6], and BART models were trained using the ‘BART’ package in R [24]. The simulation environment was coded in R v. 4.2.1. For continuous outcomes, XGBoost models were configured to minimize the squared error loss function with root mean squared error (RMSE) as the evaluation metric, while for binary outcomes, models employed logistic regression with the deviance as the evaluation metric. Hyperparameter tuning was performed via 5-fold cross-validation over a predefined grid to mitigate overfitting, adjusting learning rate (0.02, 0.05, 0.1, 0.3), maximum tree depth (3, 5, 7), and subsample ratio of features (0.8, 1.0). BART models were fitted using default hyperparameters, including a sum of 200 trees, a base prior of 0.95, and a power prior of 2, chosen to penalize excessive branch growth (see Section A2 in the Supplemental Material). Each model was trained using a single chain MCMC procedure, consisting of 500 burn-in iterations followed by 100 retained iterations for analysis.
@@ 4.3 Simulation Results
Tables 1 and S7 present the detailed results for binary and continuous outcomes. Regarding subgroup-specific treatment effect estimation, the results indicate that the “naïve” method with BART slightly outperforms the “naïve” method with XGBoost in terms of bias, variance, and MSE. Both XGBoost-trained and BART-trained models exhibit bias toward the null in subgroup-specific treatment effect estimation, though the bias in BART is slightly smaller than that in XGBoost. Additionally, the variance of the BART-trained model across 200 simulation runs is lower, indicating greater stability. As a result, the MSE is smaller for BART than for XGBoost. However, the average standard error (SE) of the ATE for both the “naïve” frequentist method and the native method with BART is smaller than that of the proposed Bayesian uncertainty-corrected method, and their 95% Confidence Interval (CI) coverages fall below 95%, indicating that the “naïve” methods underestimate the uncertainty in their estimates. In contrast, the two-stage Bayesian method maintains strong precision and accuracy while providing coverage that is close to the nominal 95%, demonstrating its superior performance in both uncertainty estimation and overall reliability.
@@ 5.1 Data Description
We apply the proposed two-stage sampling-based approach to evaluating treatment effects to data from $n=2341$ patients in the COMPILE COVID-19 clinical trial. The study evaluates the effectiveness of the COVID-19 convalescent plasma (CCP) treatment for hospitalized COVID-19 patients not on mechanical ventilation at the time of randomization [27]. The primary outcome is a binary outcome derived from the World Health Organization (WHO) 11-point clinical scale, which assesses COVID-19 severity with values ranging from 0 (no infection) to 10 (death) [20]. (see Section C in the Supplemental Material.) Specifically, we consider four binary outcomes: (1) ventilation or death, defined as an indicator for a WHO score $\geq 7$ measured at 14 ± 1 days post-randomization (the primary outcome); (2) hospitalization, ventilation, or death, defined as an indicator for a WHO score $\geq 4$ measured at 28 ± 2 days post-randomization (secondary outcome); (3) ventilation or death, defined as an indicator for a WHO score $\geq 7$ measured at 28 ± 2 days post-randomization (secondary outcome); and (4) death, defined as an indicator for a WHO score of 10 at 28 ± 2 days post-randomization (secondary outcome). Park et al.[21] developed a Treatment Benefit Index to identify hospitalized COVID-19 patients who may benefit from CCP, however, their method ignores uncertainty in estimating data-driven model-based benefit groups. The treatment variable is a binary indicator (1 for CCP treatment, 0 for control). Baseline covariates used in our study are as follows:
@@ 5.2 Application Results
To estimate the estimand in each subgroup, we implemented a Monte Carlo (MC) cross-validation (CV) procedure with 100 repetitions, as in Goligher et al. [11]. MC–CV is employed in the application to mitigate the overfitting risk inherent to single-sample analyses, whereas the simulation study uses genuinely independent design and evaluation samples and therefore does not require cross-validation. In each MC iteration, we followed the following steps:
@@
A final predicted estimand was obtained by taking the median point estimate and confidence interval across 100 MC replications [11]. The reported metrics in Table S10 are as follow: 1) OR: exponentiated median of the log ORs; 2) $\mbox{SE}_{logOR}$: median of the standard errors of the log ORs; and 3) 95% CI; exponentiated median of the lower and upper bounds of the 95% confidence intervals of the log ORs.
@@ 6 Discussion
Beyond statistical considerations, the LR framework can have clinical and regulatory value. In trials with null or inconclusive average effects, LR analyses may suggest that a treatment benefits a subset of patients; in trials with positive averages, they can help characterize where responses are stronger. Within this framework, our Bayesian approach explicitly propagates uncertainty from subgroup identification to subgroup-specific inference, reducing the risk of overconfident conclusions. LR analyses also connect to enrichment in trial design. U.S. FDA guidance describes enrichment as prospectively selecting patients more likely to respond than the broader target population, thereby improving power [28]. LR subgroups identified within an RCT offer a data-driven input to such enrichment, because potential responders are characterized using baseline features. We note that LR subgrouping based on prognosis under treatment $s(x)=E[Y\mid T=1,x]$ does not, in general, guarantee enrichment for individuals with the largest incremental treatment advantage $\Delta(x)=E[Y\mid T=1,x]-E[Y\mid T=0,x]$; however, even when the alignment between $s(x)$ and $\Delta(x)$ is weak or negative, our two-stage procedure continues to provide well-calibrated uncertainty quantification for treatment effects within the resulting clinically defined (rule-based) subgroup (see Supplementary Section B3).
@@
Although the method is developed in the context of an RCT, it can be extended to more general settings. In early-phase or pilot studies where only single-arm data are available, the “design” stage can rely on treated patients alone to develop a biomarker-based score, which can then be used to define subgroups in a subsequent, independent trial. The approach is also applicable to other model-based, data-driven subgroup analyses, such as those based on stratification by heterogeneous treatment effects [26, 18].
@@ A. Introduction of Tree-based Models
The XGBoost method [5] is based on the concept of gradient boosting , a machine learning technique that enhances weak learners into strong ones. Friedman et al.[10] later adapted boosting to a forward stagewise additive model that minimizes a loss function and brought forth a highly adaptable algorithm, gradient boosting machines (GBM). Specifically, a GBM uses $m$ additive functions to predict the outcome,
@@
XGBoost further mitigates overfitting using shrinkage and column subsampling. Shrinkage scales newly added weights after each boosting step, similar to traditional gradient tree boosting. Column subsampling, inspired by the Random Forest algorithm [3], selects a random subset of predictors for splitting at each step, enhancing model robustness.[5]
@@
The BART model is a Bayesian probability model, requiring prior distribution specification. Chipman et al. [7] simplified this step using independence assumptions:
@@
For $T_{j}$, the prior probability that nodes at depth $d$ are non-terminal is set as $\alpha(1+d)^{-\beta}$ with base $\alpha=0.95$ and power $\beta=2$ [7], encouraging smaller trees. Under this distribution, tree size of $\{1,2,3,4,5\}$ has a prior probability of $\{0.05,0.55,0.28,0.09,0.03\}$, respectively.
@@
Given the prior and the observed data, posterior sampling in BART models is carried out via Markov chain Monte Carlo (MCMC) based on "Bayesian back-fitting" [13], which is an application of the "simple" Gibbs sampler. Specifically, the $j$th tree $(T_{j},M_{j})$ is fit iteratively, holding all other m1m-1 trees constant. For each iteration, given $\{T_{j}\}$ and $\{M_{j}\}$ and based on the model residual $R_{i}=Y_{i}-\sum_{j=1}^{m}g(x_{i},T_{j},M_{j})$, we draw a sample from the posterior of $\sigma^{2}$ ("Step 1"). Then, we define the $j$th partial residual response